# Electromagnetic characteristics as probes into the inner structures of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states

### v1

@@ I Introduction
In previous theoretical studies, the experimental observation of the hidden-charm molecular pentaquark candidates alongside the double-charm molecular tetraquark candidate has motivated extensive investigations into the mass spectra of various double-charm molecular pentaquark states Wang:2023mdj ; Wang:2022aga ; Dong:2021bvy ; Chen:2021kad ; Yalikun:2023waw ; Wang:2023ael ; Wang:2023aob ; Sheng:2024hkf ; Wang:2025hhx ; Wang:2023eng ; Xu:2025mhc ; Yang:2024okq ; Wang:2024brl ; Duan:2024uuf ; Liu:2023clr ; Shen:2022zvd . Among them, a systematic analysis of the mass spectra of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}-$type double-charm hidden-strangeness molecular pentaquarks was conducted within the framework of the one-boson-exchange (OBE) model in Ref. Yalikun:2023waw . Their results indicate the presence of several loosely bound states with the spin-parity assignments $J^{P}=1/2^{-}$, $3/2^{-}$, and $5/2^{-}$, which can be regarded as promising double-charm hidden-strangeness molecular pentaquark candidates.
@@ II The magnetic moment properties of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecules
In Ref. Yalikun:2023waw , the mass spectra of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}-$type double-charm hidden-strangeness molecular pentaquark candidates with $J^{P}=1/2^{-}$, $3/2^{-}$, and $5/2^{-}$ were systematically investigated within the OBE model, incorporating exchanges of the light mesons $\sigma$, $\eta$, and $\phi$. As reported in previous work Yalikun:2023waw , the ten near-threshold loosely bound states $\Xi_{c}D_{s}(1/2^{-})$, $\Xi_{c}^{{}^{\prime}}D_{s}(1/2^{-})$, $\Xi_{c}D^{*}_{s}(1/2^{-})$, $\Xi_{c}D^{*}_{s}(3/2^{-})$, $\Xi_{c}^{*}D_{s}(3/2^{-})$, $\Xi_{c}^{{}^{\prime}}D^{*}_{s}(1/2^{-})$, $\Xi_{c}^{{}^{\prime}}D^{*}_{s}(3/2^{-})$, $\Xi_{c}^{*}D^{*}_{s}(1/2^{-})$, $\Xi_{c}^{*}D^{*}_{s}(3/2^{-})$, and $\Xi_{c}^{*}D^{*}_{s}(5/2^{-})$ were identified in single-channel analysis with $S$-$D$ wave mixing effects, using a phenomenologically reasonable cutoff parameter from 1.0 to 2.5 GeV. A subsequent coupled-channel analysis further corroborated the viability of several $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ states as promising double-charm hidden-strangeness molecular pentaquark candidates.
@@
Figure 1: The spatial wave functions $u(r)$ of the predicted ten $S$-wave $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks obtained from single-channel analysis at three representative binding energies of 0.5-0.5, 6.0-6.0, and 12.0-12.0 MeV.
@@
Given the consistency between our calculated mass spectra for the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}-$type double-charm hidden-strangeness molecular pentaquarks and those reported in Ref. Yalikun:2023waw , we do not reproduce the detailed mass spectrum results in this work. Instead, we focus on their spatial wave functions obtained from these mass spectrum calculations, which play essential role in determining their electromagnetic properties. It is important to note that the spatial wave functions of the hadronic molecular states depend sensitively on the input binding energies Chen:2017xat . However, these predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks have not yet been experimentally observed, it is theoretically well established that the molecular states are the loosely bound states, characterized by small binding energies and spatially extended wave functions Chen:2016qju . To systematically analyze their electromagnetic properties, we adopt three representative binding energies: 0.5-0.5, 6.0-6.0, and 12.0-12.0 MeV. As an illustration, Fig. 1 shows the spatial wave functions of the ten predicted $S$-wave $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks from single-channel analysis, corresponding to these three binding energies. As can be observed from Fig. 1, a smaller binding energy of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks corresponds to a more spatially extended wave function. This behavior is consistent with the theoretical picture of the hadronic molecular states, where the weakly bound states are expected to exhibit broader spatial distributions Chen:2016qju .
@@
Here, the notation $|^{2S+1}L_{J}\rangle$ labels the channels in the discussed system, where $S$, $L$, and $J$ represent the spin, orbital, and total angular momentum quantum numbers, respectively. The spin–orbital wave function for a channel $|^{2S+1}L_{J}\rangle$ is constructed as
@@ II.1 Framework for calculating the hadronic magnetic moment in the constituent quark model
$\mu_{[AB]}=\left\langle\Psi_{J_{[AB]},\,J_{[AB]}}\left|\sum\limits_{j}\hat{\mu}_{zj}^{\text{spin}}+\hat{\mu}_{z}^{\text{orbital}}\right|\Psi_{J_{[AB]},\,J_{[AB]}}\right\rangle.$ (2.2)
Here, $|\Psi_{J_{[AB]},\,J_{[AB]}}\rangle$ denotes the total wave function of the discussed molecular state $[AB]$, constructed as the direct product of the color, flavor, spin-orbit, and spatial components corresponding to its internal degrees of freedom. The $z$-component for both the spin magnetic moment operator $\hat{\mu}_{zj}^{\text{spin}}$ and the orbital magnetic moment operator $\hat{\mu}_{z}^{\text{orbital}}$ are defined as Liu:2003ab ; Huang:2004tn ; Zhu:2004xa ; Wang:2016dzu ; Li:2021ryu ; Zhou:2022gra ; Wang:2022tib ; Gao:2021hmv ; Wang:2023bek ; Lai:2024jfe ; Mutuk:2024ltc ; Mutuk:2024elj ; Li:2024wxr ; Lei:2024geu ; Lei:2023ttd ; Zhang:2025ame ; Wang:2022nqs ; Mutuk:2024jxf ; Guo:2023fih :
@@
As defined in Eq. (2.3), the constituent quark masses are essential inputs for calculating the hadronic magnetic moments within the constituent quark model. In this work, we adopt the following set of constituent quark masses: $m_{u}=336.00$ MeV, $m_{d}=336.00$ MeV, $m_{s}=450.00$ MeV, and $m_{c}=1680.00$ MeV Kumar:2005ei . This parameter set has not only successfully reproduced the magnetic moments of the octet and decuplet baryons where experimental data exist Schlumpf:1993rm ; Kumar:2005ei ; Ramalho:2009gk , but has also been widely employed in recent systematic studies of the magnetic moments for the hadronic molecular states Sheng:2024hkf ; Li:2021ryu ; Wang:2023bek ; Gao:2021hmv ; Wang:2022tib ; Wang:2023aob ; Wang:2023ael ; Wang:2022nqs ; Zhou:2022gra ; Lai:2024jfe ; Zhang:2025ame ; Wang:2024sbw . When accounting for SSDD wave mixing effects, the masses of the constituent hadrons within the molecular state also become crucial, as indicated in Eq. (2.4). In our numerical analysis, we use $m_{\Xi_{c}^{+}}=2467.71$ MeV, $m_{\Xi_{c}^{0}}=2470.44$ MeV, $m_{\Xi_{c}^{\prime+}}=2578.20$ MeV, $m_{\Xi_{c}^{\prime 0}}=2578.70$ MeV, $m_{\Xi_{c}^{*+}}=2645.10$ MeV, $m_{\Xi_{c}^{*0}}=2646.16$ MeV, $m_{D_{s}^{+}}=1968.35$ MeV, and $m_{D_{s}^{*+}}=2112.20$ MeV for the baryons and mesons masses, taking the values from the Particle Data Group ParticleDataGroup:2024cfk .
@@
where $I$ and $I_{z}$ represent the total isospin and its $z$-component, respectively. And the spin wave functions $|S,S_{z}\rangle$ for the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ systems are constructed via the following coupling scheme:
@@
Here, $S_{A}$ denotes the $z$-component of the spin for the corresponding hadron $A$, and $C^{S,S_{z}}_{s_{1},m_{1};s_{2},m_{2}}$ represents the Clebsch–Gordan coefficient in the coupling scheme.
@@ II.2.2 $S$-$D$ wave mixing analysis
Figure 3: The magnetic moments $\mu_{[AB]}$ of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states in SSDD wave mixing analysis. The corresponding single-channel results are shown as the horizontal lines for comparison.
When studying the magnetic moments of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks with SSDD wave mixing effects included, the spatial wave functions of both the $S$- and $D$-wave components are required as essential inputs. For this purpose, we adopt three typical binding energies of 0.5-0.5, 6.0-6.0, and 12.0-12.0 MeV to systematically examine their resulting magnetic moments under SSDD wave mixing analysis. The corresponding values for the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks are summarized in Fig. 3.
Through a systematic comparison of the magnetic moments of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks obtained from single-channel analysis and SSDD wave mixing analysis, we observe that the inclusion of SSDD wave mixing effects leads to only minor modifications. Moreover, the resulting magnetic moments show negligible dependence on the chosen binding energies. This behavior can be attributed to the very small $D$-wave component in the total spatial wave function, which leads to a correspondingly minor contribution to the overall magnetic moment.

### v2

@@ I Introduction
In previous theoretical studies, the experimental observation of the hidden-charm molecular pentaquark candidates alongside the double-charm molecular tetraquark candidate has motivated extensive investigations into the mass spectra of various double-charm molecular pentaquark states Wang:2023mdj ; Wang:2022aga ; Dong:2021bvy ; Chen:2021kad ; Yalikun:2023waw ; Wang:2023ael ; Wang:2023aob ; Sheng:2024hkf ; Wang:2025hhx ; Wang:2023eng ; Xu:2025mhc ; Yang:2024okq ; Wang:2024brl ; Duan:2024uuf ; Liu:2023clr ; Shen:2022zvd . Among them, a systematic analysis of the mass spectra of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}-$type double-charm hidden-strangeness molecular pentaquarks was conducted within the framework of the one-boson-exchange (OBE) model in Ref. Yalikun:2023waw . Their results indicate the presence of several loosely bound states with $J^{P}={1/2}^{-}$, ${3/2}^{-}$, and ${5/2}^{-}$, which can be regarded as promising double-charm hidden-strangeness molecular pentaquark candidates.
@@ II The magnetic moment properties of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecules
In Ref. Yalikun:2023waw , the mass spectra of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}-$type double-charm hidden-strangeness molecular pentaquark candidates with $J^{P}={1/2}^{-}$, ${3/2}^{-}$, and ${5/2}^{-}$ were systematically investigated within the OBE model, incorporating exchanges of the light mesons $\sigma$, $\eta$, and $\phi$. As reported in previous work Yalikun:2023waw , the ten near-threshold loosely bound states $\Xi_{c}D_{s}({1/2}^{-})$, $\Xi_{c}^{{}^{\prime}}D_{s}({1/2}^{-})$, $\Xi_{c}D^{*}_{s}({1/2}^{-})$, $\Xi_{c}D^{*}_{s}({3/2}^{-})$, $\Xi_{c}^{*}D_{s}({3/2}^{-})$, $\Xi_{c}^{{}^{\prime}}D^{*}_{s}({1/2}^{-})$, $\Xi_{c}^{{}^{\prime}}D^{*}_{s}({3/2}^{-})$, $\Xi_{c}^{*}D^{*}_{s}({1/2}^{-})$, $\Xi_{c}^{*}D^{*}_{s}({3/2}^{-})$, and $\Xi_{c}^{*}D^{*}_{s}({5/2}^{-})$ were identified in single-channel analysis with $S$-$D$ wave mixing effects, using a phenomenologically reasonable cutoff parameter from 1.0 to 2.5 GeV. A subsequent coupled-channel analysis further corroborated the viability of several $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ states as promising double-charm hidden-strangeness molecular pentaquark candidates. Given the consistency between our calculated mass spectra for the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}-$type double-charm hidden-strangeness molecular pentaquarks and those reported in Ref. Yalikun:2023waw , we do not reproduce the detailed mass spectrum results in this work.
@@
Figure 1: The spatial wave functions $u(r)$ of ten predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks obtained from the single-channel analysis at three representative binding energies of 0.5-0.5, 6.0-6.0, and 12.0-12.0 MeV.
@@
However, the binding energies of these predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states remain ambiguous, due to the lack of experimental observation of such states. Within the constituent quark model, the determination of the electromagnetic properties of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states relies on the spatial wave functions and the masses of these molecular states. Notably, both the spatial wave functions and the masses are closely related to their binding energies Chen:2017xat . Although the exact binding energies of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states are unknown, it is theoretically well established that such molecular states are loosely bound Chen:2016qju , characterized by small binding energies and spatially extended wave functions. In this study, we consider three representative binding energies 0.5-0.5, 6.0-6.0, and 12.0-12.0 MeV to investigate the electromagnetic properties of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states. In the following, we discuss their spatial wave functions. As an illustration, Fig. 1 shows the spatial wave functions $u(r)$ of ten predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks corresponding to these three binding energies, which can be obtained from the quantitative study of their mass spectrum. As can be observed from Fig. 1, a smaller binding energy of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks corresponds to a more spatially extended wave function. This behavior is consistent with the theoretical picture of the hadronic molecular states, where the weakly bound states are expected to exhibit broader spatial distributions Chen:2016qju .
@@
Here, the notation $|^{2S+1}L_{J}\rangle$ labels the channels in the discussed system, where $S$, $L$, and $J$ represent the spin, orbital, and total angular momentum quantum numbers, respectively. The spin-orbital wave function for a channel $|^{2S+1}L_{J}\rangle$ is constructed as
@@ II.1 Framework for calculating the hadronic magnetic moment in the constituent quark model
$\mu_{[AB]}=\left\langle J_{[AB]},\,J_{[AB]}\left|\sum\limits_{j}\hat{\mu}_{zj}^{\text{spin}}+\hat{\mu}_{z}^{\text{orbital}}\right|J_{[AB]},\,J_{[AB]}\right\rangle.$ (2.2)
Here, $|J_{[AB]},\,J_{[AB]}\rangle$ denotes the total wave function of the discussed molecular state $[AB]$, constructed as the direct product of the color, flavor, spin-orbit, and spatial components corresponding to its internal degrees of freedom. The $z$-component for both the spin magnetic moment operator $\hat{\mu}_{zj}^{\text{spin}}$ and the orbital magnetic moment operator $\hat{\mu}_{z}^{\text{orbital}}$ are defined as Liu:2003ab ; Huang:2004tn ; Zhu:2004xa ; Wang:2016dzu ; Li:2021ryu ; Zhou:2022gra ; Wang:2022tib ; Gao:2021hmv ; Wang:2023bek ; Lai:2024jfe ; Mutuk:2024ltc ; Mutuk:2024elj ; Li:2024wxr ; Lei:2024geu ; Lei:2023ttd ; Zhang:2025ame ; Wang:2022nqs ; Mutuk:2024jxf ; Guo:2023fih :
@@
As defined in Eq. (2.3), the constituent quark masses are essential inputs for calculating the hadronic magnetic moments within the constituent quark model. In this work, we adopt the following set of constituent quark masses: $m_{u}=336.00$ MeV, $m_{d}=336.00$ MeV, $m_{s}=450.00$ MeV, and $m_{c}=1680.00$ MeV Kumar:2005ei . This parameter set has not only successfully reproduced the magnetic moments of the octet and decuplet baryons where experimental data exist Schlumpf:1993rm ; Kumar:2005ei ; Ramalho:2009gk , but has also been widely employed in recent systematic studies of the magnetic moments for the hadronic molecular states Sheng:2024hkf ; Li:2021ryu ; Wang:2023bek ; Gao:2021hmv ; Wang:2022tib ; Wang:2023aob ; Wang:2023ael ; Wang:2022nqs ; Zhou:2022gra ; Lai:2024jfe ; Zhang:2025ame ; Wang:2024sbw . When accounting for $S$-$D$ wave mixing effects, the masses of the constituent hadrons within the molecular state also become crucial, as indicated in Eq. (2.4). In our numerical analysis, we use $m_{\Xi_{c}^{+}}=2467.71$ MeV, $m_{\Xi_{c}^{0}}=2470.44$ MeV, $m_{\Xi_{c}^{\prime+}}=2578.20$ MeV, $m_{\Xi_{c}^{\prime 0}}=2578.70$ MeV, $m_{\Xi_{c}^{*+}}=2645.10$ MeV, $m_{\Xi_{c}^{*0}}=2646.16$ MeV, $m_{D_{s}^{+}}=1968.35$ MeV, and $m_{D_{s}^{*+}}=2112.20$ MeV for the baryons and mesons masses, taking the values from the Particle Data Group ParticleDataGroup:2024cfk .
@@ II.2.1 Single-channel analysis
where II and $I_{z}$ represent the total isospin and its $z$-component, respectively. And the spin wave functions $|S,S_{z}\rangle$ for the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ systems are constructed via the following coupling scheme:
@@
Here, $S_{A}$ denotes the $z$-component of the spin for the corresponding hadron $A$, and $C^{S,S_{z}}_{s_{1},m_{1};s_{2},m_{2}}$ represents the Clebsch-Gordan coefficient in the coupling scheme.
@@
As important input parameters, the constituent quark masses play a crucial role in determining the electromagnetic properties of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states within the constituent quark model. However, due to the lack of relevant experimental data, these masses cannot be determined with high precision. These parameter set adopted in this work has not only successfully reproduced the magnetic moments of the octet and decuplet baryons, for which experimental data are available, but has also been widely employed in recent systematic studies of the electromagnetic properties of the hadronic molecular states. To assess the impact of parametric uncertainties, we vary all constituent quark masses within a physically motivated range of approximately $\pm 10\%$ when discussing the magnetic moments of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states in single-channel analysis. This range covers the typical variations found in different constituent quark masses parametrization in Ref. Majethiya:2009vx . The resulting ranges for the magnetic moments will be indicated as superscripts and subscripts attached to the central values (see Table 1). From Table 1, we find that while the constituent quark masses affect the magnetic moments of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states, their influence is not significant.
Table 1: The magnetic moments of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states in single-channel analysis, in units of $\mu_{N}$. The central values are calculated using the fixed constituent quark masses, while the superscripts and subscripts denote the upper and lower bounds obtained from varying the constituent quark masses within approximately $\pm 10\%$.
Molecules $I_{z}=+\frac{1}{2}$ $I_{z}=-\frac{1}{2}$
$\Xi_{c}D_{s}]|\frac{1}{2}^{-}\rangle$ $0.37^{+0.04}_{-0.03}$ $0.37^{+0.04}_{-0.03}$
$\Xi_{c}^{{}^{\prime}}D_{s}]|\frac{1}{2}^{-}\rangle$ $0.65^{+0.19}_{-0.18}$ $1.21^{+0.11}_{-0.13}$
$\Xi_{c}D_{s}^{*}]|\frac{1}{2}^{-}\rangle$ $0.59^{+0.09}_{-0.08}$ $0.59^{+0.09}_{-0.08}$
$\Xi_{c}D_{s}^{*}]|\frac{3}{2}^{-}\rangle$ $1.44^{+0.16}_{-0.13}$ $1.44^{+0.16}_{-0.13}$
$\Xi_{c}^{*}D_{s}]|\frac{3}{2}^{-}\rangle$ $1.54^{+0.31}_{-0.28}$ $1.25^{+0.19}_{-0.21}$
$\Xi_{c}^{{}^{\prime}}D_{s}^{*}]|\frac{1}{2}^{-}\rangle$ $0.49^{+0.14}_{-0.13}$ $1.11^{+0.12}_{-0.10}$
$\Xi_{c}^{{}^{\prime}}D_{s}^{*}]|\frac{3}{2}^{-}\rangle$ $1.72^{+0.31}_{-0.28}$ $0.14^{+0.23}_{-0.23}$
$\Xi_{c}^{*}D_{s}^{*}]|\frac{1}{2}^{-}\rangle$ $0.50^{+0.21}_{-0.20}$ $1.05^{+0.14}_{-0.16}$
$\Xi_{c}^{*}D_{s}^{*}]|\frac{3}{2}^{-}\rangle$ $1.56^{+0.28}_{-0.24}$ $0.49^{+0.19}_{-0.20}$
$\Xi_{c}^{*}D_{s}^{*}]|\frac{5}{2}^{-}\rangle$ $2.61^{+0.43}_{-0.38}$ $0.19^{+0.31}_{-0.31}$
@@ II.2.2 $S$-$D$ wave mixing analysis
Figure 3: The magnetic moments $\mu_{[AB]}$ of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states in $S$-$D$ wave mixing analysis. The corresponding single-channel results are shown as the horizontal lines for comparison.
When studying the magnetic moments of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks with $S$-$D$ wave mixing effects included, the spatial wave functions of both the $S$- and $D$-wave components are required as essential inputs. For this purpose, we adopt three typical binding energies of 0.5-0.5, 6.0-6.0, and 12.0-12.0 MeV to systematically examine their resulting magnetic moments under $S$-$D$ wave mixing analysis. The corresponding values for the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks are summarized in Fig. 3.
Through a systematic comparison of the magnetic moments of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks obtained from single-channel analysis and $S$-$D$ wave mixing analysis, we observe that the inclusion of $S$-$D$ wave mixing effects leads to only minor modifications. Moreover, the resulting magnetic moments show negligible dependence on the chosen binding energies. This behavior can be attributed to the very small $D$-wave component in the total spatial wave function, which leads to a correspondingly minor contribution to the overall magnetic moment.